OpenStudy (anonymous):

What is the equation of the line passing through (2, –2) and perpendicular to the line y = –x + 3 in slope-intercept form?

OpenStudy (anonymous):

you have k=-1 for prependicular lines you have \[\LARGE k\cdot k_1=-1 \] \[\LARGE-1\cdot k_1=-1 \] \[\LARGE - k_1=-1 \] \[\LARGE k_1=1 \] so plug it into... \[\LARGE y-y_1=k(x-x_1 )\]

OpenStudy (anonymous):

sorry... I didn't mentioned where k=-1 came from.. \[\LARGE y=-x+3\] \[\Huge y=\underbrace{\boxed{-1}}_{k}\cdot x+3\]